OpenStudy (anonymous):

A catapult launches a boulder with an upward velocity of 112 ft/s. The height of the boulder, h, in feet after t seconds is given by the function h= -16t^2 + 112t + 30 . How long does it take the boulder to reach its maximum height? What is the boulder’s maximum height? Round to the nearest hundredth, if necessary.

OpenStudy (anonymous):

A. 7 s; 30 ft B. 3.5 s; 366 ft C. 3.5 s; 618 ft D. 3.5 s; 226 ft

OpenStudy (ksaimouli):

is this precal

OpenStudy (baldymcgee6):

use the vertex formula

OpenStudy (anonymous):

@baldymcgee6 how do I do that

OpenStudy (baldymcgee6):

-b/2a

OpenStudy (ksaimouli):

can u use calculator

OpenStudy (ksaimouli):

D

OpenStudy (ksaimouli):

D. 3.5 s; 226 ft

OpenStudy (ksaimouli):

i am sure

OpenStudy (anonymous):

WIth the calculator what would I divide or Minus?

OpenStudy (baldymcgee6):

is this for a calculus class or not?

OpenStudy (ksaimouli):

i think it is not @baldymcgee6

OpenStudy (baldymcgee6):

you don't know what class you're in? haha

OpenStudy (anonymous):

Its Algebra 2

OpenStudy (ksaimouli):

r u adressing me

OpenStudy (ksaimouli):

use calculator

OpenStudy (ksaimouli):

find max point

OpenStudy (anonymous):

How @ksaimouli, i wrote everything into the calculator and I got -114 that is not correct answer

OpenStudy (baldymcgee6):

-b/2a = -112/2(16) = 3.5 s t= 3.5s h = -16t^2 + 112t + 30 h = -16(3.5)^2 +122(3.5) +30 h = 226

OpenStudy (ksaimouli):

adjust the windows to y=300

OpenStudy (baldymcgee6):

@chicagochica5 i told you how, use the vertex formula, and this will give you the time at which it is at its maximum

OpenStudy (ksaimouli):

say 2nd and calc and press find max

OpenStudy (anonymous):

@ksaimouli, Just I have a phone calculator

OpenStudy (ksaimouli):

@baldymcgee6 is that formula used in calculus

OpenStudy (baldymcgee6):

@chicagochica5 it is better to do it by hand, @ksaimouli is suggesting you use a graphing calculator, which is totally possible and easier, but it doesn't help you learn and sometimes you are not allowed a calculator.

OpenStudy (ksaimouli):

no u should have ti-83

OpenStudy (ksaimouli):

yup u r right

OpenStudy (baldymcgee6):

@ksaimouli some classes do not allow a graphing calculator such as the ti 83, and no it definitely not a calculus formula.

OpenStudy (ksaimouli):

but i confuses when we say another and their teacher says totally different

OpenStudy (baldymcgee6):

its simply a parabolic formula to find the vertex of a parabola.

OpenStudy (anonymous):

Thanks @ksaimouli and @baldymcgee6

OpenStudy (ksaimouli):

welcome

OpenStudy (baldymcgee6):

you are welcome @chicagochica5